Let us denote Sn = an + bn + cn for arbitrary numbers a, b, c. 

It is known that S1 = 8, S2 = 66, S3 = 536 for some values of

a, b, c. What is the smallest possible value of S242 — S41 S43?


1
Expert's answer
2020-11-30T21:03:23-0500

"S_{42}^2 - S_{41}*S_{43} = (a^{42} + b^{42} +c^{42})^2 -\\
(a^{41} + b^{41} +c^{41})*(a^{43} + b^{43} +c^{43}) = \\
 a^{84} + b^{84} +c^{84} +2a^{42}b^{42} + 2a^{42}c^{42} +\\
+2b^{42}c^{42} -a^{84} - b^{84} -c^{84} - a^{41}(b^{43} +c^{43})-\\
-b^{41}(a^{43} +c^{43})-c^{41}(a^{43} +b^{43}) = \\
=2a^{42}b^{42} + 2a^{42}c^{42} +2b^{42}c^{42} - a^{41}(b^{43} +c^{43})-\\
-b^{41}(a^{43} +c^{43})-c^{41}(a^{43} +b^{43}) \\
\begin{cases}
a+b+c = 8\\
a^2+b^2+c^2 = 66\\
a^3+b^3+c^3 = 536
\end{cases}\\
\text{since the system is symmetric there will be only one solutions}:\\
\begin{cases}
c = 8 -a- b\\
a^2+b^2+c^2 = 66\\
a^3+b^3+c^3 = 536
\end{cases}\\
\begin{cases}
a^2+b^2+64+a^2+b^2-16a-16b-2ab = 66\\
a^3+b^3+c^3 = 536
\end{cases}\\
\begin{cases}
2a^2+2b^2-16a-16b-2ab = 2\\
a^3+b^3+512 - 192(a+b) + 24(a+b)^2-(a+b)^3 = 536
\end{cases}\\
a = 0 , b= 4 -\sqrt{17} , c=4 +\sqrt{17}\\
2(4 -\sqrt{17})^{42}(4 +\sqrt{17})^{42} - \\
-(4 -\sqrt{17})^{41}(4 +\sqrt{17})^{43} - \\
-(4 +\sqrt{17})^{41}(4 -\sqrt{17})^{43} = \\
= -(4 -\sqrt{17})^{41}(4 +\sqrt{17})^{41}*\\
*((4 -\sqrt{17})^{2} +2(4 -\sqrt{17})(4 +\sqrt{17}) + (4 +\sqrt{17})^{2}) = \\
= - (16-17)^{41} * (64) = -(-1)^{41} *(64) = 64\\
answer: 64"